# Uncovering Bias Paths with LLM-guided Causal Discovery: An Active Learning and Dynamic Scoring Approach

### v1

@@ 6 Datasets
To evaluate the performance of the proposed LLM-based CD framework with AL and dynamic scoring, a series of experiments are conducted on benchmark networks. These networks are described in the following.
@@
The Child network is a medium-scale Bayesian network simulating causal relationships relevant to congenital heart disease in newborns (Spiegelhalter et al. 1993). It includes 20 nodes and 25 edges spanning medical, parental, and environmental factors. Variables include both categorical and continuous types, covering clinical features such as birth asphyxia, lung function, cardiac mix, chest X-ray results, and outcomes like mortality and hospital stay.
Neuropathic Pain Causal Network.
The Neuropathic Pain Diagnosis Simulator (Tu et al. 2019) models causal relationships underlying neuropathic pain using a graph of 221 nodes and 770 edges, based on real patient data. It includes symptom, pattern, and pathophysiological diagnoses from actual patient data.
As these benchmarks lacks sensitive variables, we do not assess fairness; evaluation focuses solely on structural accuracy.
@@ 7 Experiments
To evaluate the proposed LLM-based CD framework, we conduct a series of experiments on the introduced datasets. The primary goal is to assess the framework’s ability to improve causal graph accuracy and recover fairness-relevant pathways under varying configurations and noise conditions. This section outlines our experimental design, evaluation metrics, and baseline comparisons.
7.1 Baseline Methods
We compare our method against the following CD approaches:
Peter-Clark (PC) Algorithm.
A constraint-based method using conditional independence tests to iteratively prune edges from a fully connected graph, followed by orienting edges to form a DAG. It assumes causal sufficiency and the faithfulness assumption, relying heavily on accurate independence tests (Spirtes and Glymour 1991).
Greedy Equivalence Search (GES) Algorithm.
A score-based method that operates in two phases: a forward phase that incrementally adds edges to improve a score function (e.g., BIC), and a backward phase that removes edges to optimize the graph structure further. It assumes causal sufficiency and focuses on equivalence classes of DAGs to improve computational efficiency (Meek 1997).
NOTEARS Algorithm.
Formulates CD as a continuous optimization problem by introducing a differentiable acyclicity constraint. This allows it to learn DAGs using gradient-based optimization, balancing data fit and sparsity (Zheng et al. 2018).
DAGMA Algorithm.
Integrates neural networks and optimization techniques to model nonlinear causal relationships. It uses a masked autoencoder architecture to enforce sparsity and incorporates a differentiable acyclicity condition to ensure that the resulting graph is a valid DAG (Bello, Aragam, and Ravikumar 2022).
LLM-based Pairwise Method.
Utilizes LLMs to infer causal relationships through pairwise queries based on metadata and variable descriptions. While intuitive, it scales poorly with graph size (Kıcıman et al. 2023).
LLM-based BFS Method.
LLM-based BFS appraoch introduced by Jiralerspong et al. (Jiralerspong et al. 2024) which was discussed in the Sections 1 and 3.
All baselines are implemented using the open-source framework from (Jiralerspong et al. 2024) for consistency.
7.2 Experimental Design
We apply Bayesian optimization with a Gaussian Process (GP) surrogate model to explore the influence of key hyperparameters on performance. This enables tuning of our query scoring and LLM prompting strategy to enhance edge discovery and fairness-related signal recovery (Mockus 1994; Snoek, Larochelle, and Adams 2012). Optimization is performed using gp_minimize from scikit-optimize.
The hyperparameters optimized include:
  • Scoring Weights: Relative contributions of MI, PCorr, and query history in Equation 1.
  • Score threshold: Minimum dynamic score for querying a variable pair.
  • LLM Temperature: Controls the stochasticity of LLM responses.
  • Max Iterations: Maximum number of AL iterations per trial.
To evaluate robustness, we divide trials into chunks that sample different regions of the hyperparameter space. Each chunk records graph outputs and performance metrics, and the top-performing configuration from each run is used for final reporting. Full details on the hyperparameters used in each set of experiments are provided in Appendix B.
We evaluate all methods on four synthetic datasets generated from the Adult-based causal graph (Section 5) using different random seeds, which preserve the underlying structure while introducing independent data realizations. Performance metrics are averaged across these seeds. Fairness pathway analyses are conducted only on the Adult-based datasets, focusing on paths from sensitive attributes (sex, race, age) to the outcome variable (income).
We also evaluate all methods on the Neuropathic and Child benchmark networks, averaging results for LLM-based methods across five independent runs.
7.3 Performance Evaluation Metrics
We assess graph reconstruction performance using standard structure-learning metrics measured against the ground-truth DAG.
Precision and Recall.
$\text{Precision}=\frac{\text{TP}}{\text{TP + FP}},\quad\text{Recall}=\frac{\text{TP}}{\text{TP + FN}}$
F1 Score.
[…]
Edge Count Comparison.
We compare the number of predicted edges with the number in the ground-truth graph.
DAG Validation.
Checks acyclicity of the predicted structure.
Normalized Hamming Distance (NHD).
$\text{NHD}=\frac{\text{Mismatches}}{n^{2}}$
where n refers to the number of nodes in the graph.
Adjacency Accuracy.
Accuracy=Correct $\text{Accuracy}=\frac{\text{Correct entries}}{n^{2}}$
@@ 8 Results and Analyses
8.1 Causal Discovery Performance Comparison Between Methods
Table 1: Performances of Methods on the Adult-based Synthetic, Child and Neuropathic Causal Networks
@@ 8.1 Causal Discovery Performance Comparison Between Methods
Table 1 summarizes performance across the synthetic Adult-based, Child, and Neuropathic networks using accuracy, precision, recall, F1 score, and NHD. Across all datasets, the proposed method consistently outperforms baseline approaches in F1 score, precision, and structural alignment with the ground-truth graph. The proposed method generally leads to conservative graph reconstruction compared to the baselines.
@@
The proposed method outperforms all baselines in most metrics except recall. It achieves the highest F1 score (0.585), accuracy (0.413), precision (0.792), and the lowest NHD (0.109) and NHD ratio (0.415). This demonstrates its superiority in structural accuracy and edge prediction while producing a more concise graph (17 predicted edges) than other methods.
LLM Pairwise method achieves the highest recall (0.813) but overpredicted edges (69). Classical methods like GES and PC achieved moderate F1 scores (0.473 and 0.382, respectively), but their graphs had higher NHDs and more structural violations. Continuous optimization methods (NOTEARS and DAGMA) underperform across most metrics, indicating sensitivity to noise and difficulty capturing fairness-relevant structure.
Child Network (20 nodes, 25 edges).
The proposed method achieves the highest F1 score (0.533), precision (0.601), and accuracy (0.364), outperforming both statistical and LLM-based baselines. It also achieves the lowest NHD ratio (0.467), indicating better alignment with the ground-truth structure. Although NOTEARS achieves a slightly better raw NHD, our method offers a better precision-recall balance. The LLM Pairwise method shows strong recall (0.559) but overpredicts edges, leading to lower precision and weaker structural alignment.
Neuropathic Network (221 nodes, 770 edges).
The proposed method achieves the highest F1 score (0.136), the highest precision (0.690), and the lowest normalized Hamming distance (NHD) ratio (0.864) across all tested methods. While these metrics are relatively low in absolute terms, they reflect a broader trend where all evaluated methods struggle to recover the full causal structure in this large, high-dimensional graph. This underscores the inherent difficulty of CD in settings characterized by sparse, noisy dependencies and complex variable interactions. Still, the proposed method exhibits higher performance compared to baseline approaches by achieving a better balance between edge prediction accuracy and false positive control.
The LLM-BFS baseline fails entirely on this dataset (F1 = 0, accuracy = 0), likely due to its static query strategy being overwhelmed by the Neuropathic graph’s high dimensionality and sparsity. This underscores the limitations of fixed traversal methods in large-scale settings. In contrast, our method’s use of dynamic scoring and AL improves query prioritization, enabling more effective discovery under complex conditions.
@@
In addition to performance, we also note qualitatively that our method scales more efficiently than optimization-based approaches (e.g., DAGMA, NOTEARS), and the LLM-pairwise method while offering greater structural fidelity. We also perform a complexity analysis described in Appendix E showing how our proposed method could theoretically scale better than the LLM BFS method on significantly large sparse networks, characteristics which real-world networks usually exhibit. This balance that the proposed method provides is evident in the Neuropathic setting, where computationally intensive baselines suffer performance degradation or scalability bottlenecks. Due to these scalability and resource limitations, experiments using the GES and LLM pairwise methods were not completed on the Neuropathic graph.
Reproducibility and Implementation Discrepancies.
We observe lower performance for LLM-BFS and LLM-Pairwise compared to results reported in Jiralerspong et al. (2024), particularly on the Child and Neuropathic networks. This may be due to (1) non-determinism and ongoing updates to models like GPT-4, which may impact consistency across evaluations, especially in semantically ambiguous cases. (2) implementation differences in token limits or message handling, (4) lack of access to original variable descriptions for the Neuropathic dataset, affecting metadata quality. Despite these differences, our implementation adheres to the BFS protocol and reveals its limitations in high-dimensional settings, further motivating our enhancements through dynamic scoring and AL.
8.2 Fairness Pathway Recovery and Bias Attribution
We assess each method’s fairness-awareness by analyzing causal paths from sex, age, and race to income in the Adult-based synthetic network. The ground-truth graph contains 2 direct and 25 indirect fairness paths, with most influence mediated (Indirect = 0.93, Direct = 0.07). A total effect (TE) of 4.89 and normalized bias (C_bias = 28.46) indicate substantial structural bias, aligning with real-world disparities (Appendix D).
We compare each method’s recovered graph in terms of these fairness-aware path metrics. A table detailing the averages of all the fairness and path-related scores for the methods across seed runs is included in Appendix D, as well as a visualization of the graph generated by the outputs of the proposed method.
Figure 2 shows the number of direct, indirect, and spurious paths per method. Due to the wide variance where some methods generate fewer than 5 paths (e.g., NOTEARS), others over 600 (e.g., GES), we use a log-scaled y-axis to preserve readability and comparability without distorting proportions.
Figure 2: Log-scaled path type counts (direct, indirect, spurious) across methods. The log scale enables fair comparison by compressing the large variation in total path counts.
Our proposed method is the only approach to correctly identify both direct fairness paths from sex and race to income, while correctly omitting a spurious direct path from age, demonstrating both sensitivity to true bias and restraint against overgeneralization.
In contrast, LLM Pairwise and BFS methods introduce an incorrect direct path from age, yielding three direct paths in total. This inflated their $C_{\text{bias}}$ values to appear closer to the baseline but was driven by over-attribution rather than accurate recovery. These methods tend to over-attribute causal influence to demographically meaningful variables due to LLM priors and lack of structural constraints, which can amplify perceived bias.
Classical statistical methods, including PC, GES, NOTEARS, and DAGMA, largely underperform. DAGMA and NOTEARS fail to identify any valid paths (direct or indirect) from the sensitive attributes to income, producing exclusively spurious paths. As a result, their $C_{\text{bias}}$ values are near zero, reflecting an inability to detect embedded social bias mechanisms, likely due to their strict dependence on data-driven associations and conservative edge inclusion thresholds.
PC identifies few fairness-relevant paths, and missed both direct edges, leading to a low bias score. This aligns with its constraint-based nature, which can be overly sensitive to noise in independence tests, particularly with sparse or weak signals. GES, while slightly better at recovering indirect effects, also fails to identify direct paths. However, due to its aggressive edge addition during the forward search phase, it captured a larger number of fairness-related pathways overall, majority of which are spurious. This explains its relatively close $C_{\text{bias}}$ score to the baseline, though the accuracy of its causal assignments remains debatable. The graphs constructed by all methods are shown in Appendix D.
Overall, our proposed approach demonstrates a comparatively strong alignment with the fairness-relevant structure of the ground truth. It recovered both true direct effects from sex and race to income, while avoiding spurious paths such as a direct effect from age, which is introduced erroneously by other methods. Although the resulting $C_{\text{bias}}$ value is lower than the baseline, this may reflect the method’s prioritization of high-confidence, high-scoring queries and its dynamic scoring mechanism that discourages over-inclusion. This more conservative behavior potentially reduces the risk of false positives in fairness-sensitive settings, though it may also underestimate total bias when indirect effects are more diffuse. In addition, our method also exhibits the most stable performance across runs, as reflected by its consistently low standard deviation on key metrics such as indirect path count, total effect, and normalized bias as seen in the table in Appendix D. This reliability is important in fairness-sensitive contexts, where reproducibility and robustness are critical.
Figure 3: Comparison of normalized bias ($C_{\text{bias}}$) across methods.
8.3 Parameter Impact Analysis on Proposed Method
We analyze the Bayesian optimization trials using a Random Forest regressor to identify which hyperparameters most influence F1 performance. Max iterations have the greatest impact, followed by score threshold and temperature (Figure 4). This suggests that deeper exploration in AL improves graph construction by reducing false negatives.
Figure 5 supports this, showing F1 scores increasing with iteration budget, up to a plateau, emphasizing the trade-off between computational cost and discovery depth.
Figure 4: Feature importance of hyperparameters in predicting F1 score, computed via Random Forest regression. max_iterations has the highest predictive influence.
Figure 5: Relationship between max_iterations and F1 score. Higher iteration budgets generally yield better causal graph construction performance.
The correlation matrix (Figure 6) shows strong negative correlations among scoring weights, reflecting their competitive role in the dynamic scoring mechanism, while only max iterations correlates positively with F1. This suggests that, for smaller networks like the Adult-based synthetic graph, exploration depth outweighs weight tuning in driving discovery performance.
Figure 6: Pearson correlation heatmap between F1 score and key hyperparameters. Notably, max iterations shows strong positive correlation with F1, while weight terms are moderately anti-correlated.
As graph size increases, parameter importance shifts. In the large Neuropathic network (221 nodes), statistical weights show the strongest positive correlation with F1 (Appendix C), suggesting that MI and PCorr are crucial for narrowing the query space in high-dimensional settings.
Conversely, LLM confidence and query history weights become less influential, likely due to semantic cues being less reliable across deeper or denser graphs. This highlights how our dynamic scoring mechanism could adapt to scale, favoring semantic guidance in small graphs and statistical grounding in larger ones.
@@ 9 Discussion
Our findings show that augmenting LLM-based CD with AL and dynamic scoring improves both structural accuracy and fairness sensitivity. The proposed method consistently outperformed baselines across the Adult-based synthetic network, the Child graph, and the large-scale Neuropathic Pain network.
The dynamic scoring mechanism enables effective trade-offs between statistical dependencies and LLM-derived semantics, adapting to graph complexity. This is especially valuable in high-dimensional settings, where statistical cues helped narrow the query space. AL further optimized discovery by prioritizing informative, non-redundant queries.
In terms of fairness, our method is the only one to recover both direct paths from sensitive attributes (sex, race) to income while avoiding false positives (e.g., a spurious link from age). This indicates strong potential for fairness audits with greater precision and reduced over-attribution.
Parameter sensitivity analysis reveals that exploration budget (e.g., max iterations) is the dominant driver of performance. While LLM signals aid early search stages, statistical cues grow in importance with graph size, highlighting the adaptive nature of our hybrid scoring framework across discovery contexts.
While our findings highlight the general potential of CD methods to inform fairness assessments, they also underscore a critical caveat where such methods, particularly when applied naively, can introduce spurious pathways or erroneous connections involving sensitive attributes. Across baseline methods, we observe frequent over or under attribution of causal influence to variables like race, sex, or age, leading to inflated or misleading estimates of structural bias. This is especially problematic in fairness contexts, where false positives (e.g., wrongly inferring a direct path from a sensitive attribute to an outcome), as indicated by the misleadingly similar C_bias scores of graphs from non-performing CD methods to the baseline, may trigger unjustified policy interventions or obscure true sources of inequity. The high variance in graph structure suggests that CD outputs should be interpreted with considerable caution. Practitioners must validate inferred relationships using domain knowledge, robustness checks, and sensitivity analyses before drawing conclusions about discrimination or bias. Without such safeguards, CD risks becoming a source of misinformation rather than a tool for mitigating bias.

### v2

@@ 5 Datasets
We evaluate our framework on three benchmark networks of varying size and realism.
Synthetic Adult-Based Network.
We construct a semi-synthetic dataset based on the UCI Adult dataset (Kohavi et al. 1996), a standard benchmark in fairness research. The graph includes 15 original variables, with causal edges informed by prior work (Kilbertus et al. 2017; Nabi and Shpitser 2018). Demographics (e.g., race, sex) influence education and occupation, which in turn affect income, both directly and via mediators such as capital gains. Age and marital status act as confounders or mediators.
To simulate structural bias, we inject direct edges from race and sex to income, alongside indirect paths (e.g., sex $\rightarrow$ education $\rightarrow$ income). These allow evaluation of direct, indirect, and total effects. Details on graph construction and data generation are in the Appendix.
@@
A 20-node, 25-edge Bayesian network modeling clinical, environmental, and parental factors in congenital heart disease (Spiegelhalter et al. 1993). Variables include both categorical and continuous types.
Neuropathic Pain Network.
A large-scale clinical graph with 221 nodes and 770 edges capturing pathophysiological and symptomatic relationships in neuropathic pain diagnoses, derived from real-world patient data (Tu et al. 2019).
Since the latter two datasets lack sensitive attributes, we focus on structural accuracy rather than fairness.
@@ 6 Experiments
We evaluate our LLM-based CD framework across the benchmark datasets, focusing on graph accuracy and fairness-relevant path recovery under varying noise and hyperparameter settings.
6.1 Baselines
We compare against the following methods:
  • PC Algorithm (Spirtes and Glymour 1991): Constraint-based method using independence tests.
  • GES (Meek 1997): Score-based approach optimizing a score function (e.g., BIC) via forward-backward search.
  • NOTEARS (Zheng et al. 2018): Continuous optimization using a differentiable acyclicity constraint.
  • DAGMA (Bello, Aragam, and Ravikumar 2022): Neural network-based CD with sparsity and acyclicity constraints.
  • LLM Pairwise (Kıcıman et al. 2023): Uses LLM to infer pairwise causality from metadata.
  • LLM BFS (Jiralerspong et al. 2024): Explores graph structure via LLM-guided breadth-first querying.
All baselines are run using the open-source code from (Jiralerspong et al. 2024).
6.2 Experimental Setup
We use Bayesian optimization (gp_minimize) with a Gaussian Process surrogate to tune key hyperparameters (Mockus 1994; Snoek, Larochelle, and Adams 2012):
  • Scoring Weights: Weights for MI, PCorr, and query history (Eq. 1)
  • Score Threshold: Minimum score required for querying a pair
  • LLM Temperature: Controls randomness in responses
  • Max Iterations: AL rounds per run
Each query uses a multi-turn GPT-4 (0125-preview) prompt with prior discoveries and variable metadata (Section 3). A directed edge $X\rightarrow Y$ is added if the reply contains <Answer>Yes</Answer> and does not create a cycle. If ambiguous, no edge is added. Confidence scores are derived from the average probability of top-5 tokens based on log-probs; defaulting to 0.5 if unavailable.
Trials are divided into chunks across hyperparameter space. We report results from the best-performing configuration for each run. Full prompt design and search ranges are detailed in the Appendix.
We evaluate all methods on four synthetic variants of the Adult-based dataset (Section 5), varying random seeds to preserve the underlying structure while introducing independent data realizations. Fairness analysis targets paths from sex, race, and age to income. Structural accuracy is also assessed on the Child and Neuropathic networks, with LLM-based methods averaged over five independent runs.
6.3 Evaluation Metrics
We assess reconstruction accuracy using standard metrics:
  • Precision & Recall: Defined as $\text{Precision}=\frac{\text{TP}}{\text{TP + FP}}$ and $\text{Recall}=\frac{\text{TP}}{\text{TP + FN}}$.
  • F1 Score: Harmonic mean of precision and recall.
  • Edge Count: Number of predicted edges vs. ground truth.
  • Acyclicity: Ensures output is a valid DAG.
  • Normalized Hamming Distance (NHD): $\text{NHD}=\frac{\text{Mismatches}}{n^{2}}$, where $n$ is the number of nodes.
  • Adjacency Accuracy: $\text{Accuracy}=\frac{\text{Correct entries}}{n^{2}}$.
@@ 7 Results and Analyses
7.1 Graph Performance Comparison
@@
Table 1: Performances of Methods on the Adult-based Synthetic, Child, and Neuropathic Causal Networks. All reported values are averaged over four random seeds. Metrics exhibit minimal variance across runs ($\pm 2\times 10^{-4}$ in F1 score).
@@
Our method achieves top scores across F1 (0.585), accuracy (0.413), precision (0.792), and NHD (0.109), recovering a compact graph (17 edges) with strong alignment to ground truth. LLM Pairwise attains high recall (0.813) but overpredicts, while GES and PC show moderate F1 but higher NHDs. Optimization-based methods struggle with fairness structure.
Child Network.
Again, our method leads in F1 (0.533), accuracy (0.364), and NHD ratio (0.467), balancing edge precision and coverage. NOTEARS achieves slightly better raw NHD but worse recall. LLM Pairwise finds many indirect effects, but at the cost of low precision.
Neuropathic Network.
Despite the challenge of 221 nodes, our method yields highest F1 (0.136) and precision (0.690), outperforming baselines. All baselines degrade sharply; LLM BFS fails entirely (F1 = 0). Dynamic scoring enables better query selection under sparsity.
@@
Our method scales better than NOTEARS, DAGMA, or LLM Pairwise while offering greater structural fidelity. Complexity analysis (detailed in Appendix) shows favorable scaling theoretically in large sparse graphs, common in real-world networks. GES and LLM Pairwise were excluded from Neuropathic due to resource limits.
Reproducibility.
LLM-BFS and Pairwise underperform vs. Jiralerspong et al. (2024), likely due to (1) GPT-4 updates and nondeterminism, (2) token limits or message handling, and (3) lack of access to original variable descriptions for the Neuropathic dataset, affecting metadata quality. Our implementation adheres to the BFS protocol and exposes its limitations under scale, motivating our enhancements.
7.2 Fairness Pathway Recovery
We evaluate fairness path recovery on the Adult-based synthetic graph, where the true structure contains 2 direct and 25 indirect paths from sex, age, and race to income. TE = 4.89 and $C_{\text{bias}}=28.46$.
Figure 2: Log-scaled path counts (direct, indirect, spurious) per method.
Our method uniquely recovers both true direct paths (from sex and race) while omitting the spurious age path, showing strong fairness alignment. LLM methods over-attribute (3 direct paths) due to LLM priors, inflating $C_{\text{bias}}$. PC and GES miss direct effects; NOTEARS and DAGMA capture none, yielding near-zero $C_{\text{bias}}$.
Figure 3: Normalized bias ($C_{\text{bias}}$) across methods.
Overall, our method prioritizes high-confidence discovery, avoiding false positives and maintaining robustness across seeds as reflected by its consistently low standard deviation on key metrics such as indirect path count, total effect, and normalized bias as seen in the table in Appendix. Its conservative nature may underestimate bias but enhances reliability in fairness-critical settings.
7.3 Ablation and Parameter Influence
Bayesian optimization trials were analyzed via Random Forest regression to assess parameter impact on F1 score.
Figure 4: Hyperparameter importance on F1 score.
Findings.
Max iterations have the strongest effect, suggesting deeper AL improves recovery. Score threshold and temperature also influence performance. Correlation analysis (Figure 5) shows anti-correlation among scoring weights, indicating competitive trade-offs. In small graphs, deeper querying matters more; in large ones (e.g., Neuropathic), MI/PCorr dominate (details in Appendix). This highlights how our dynamic scoring mechanism could adapt to scale, favoring semantic guidance in small graphs and statistical grounding in larger ones.
Figure 5: F1 vs. hyperparameter correlation (Adult).
@@ 8 Discussion
Our findings demonstrate that combining LLM-based CD with AL and dynamic scoring improves both structural accuracy and fairness sensitivity across diverse datasets. The proposed method consistently outperforms statistical, LLM-only, and hybrid baselines on both small and large networks.
Dynamic scoring balances LLM semantic priors with empirical signals, adapting to graph complexity and data quality. In early iterations, LLM judgments guide exploration when statistical signals are weak. As more information is gathered, mutual information and partial correlation become dominant, reflecting a shift from semantic exploration to empirical refinement. AL enhances efficiency by prioritizing uncertain, low-redundancy variable pairs, avoiding the limitations of fixed-order strategies.
Our method uniquely recovers fairness-critical paths, e.g., sex $\rightarrow$ education $\rightarrow$ income—while avoiding spurious ones like age $\rightarrow$ income, often introduced by LLM-only baselines. This improves fairness diagnostics for impact assessments in domains like hiring or lending. While our fairness evaluation relies on a semi-synthetic benchmark, this is necessary as real-world datasets lack ground-truth causal graphs for sensitive attributes and outcomes.
However, our results also underscore broader concerns: CD methods, including ours, may introduce spurious or incomplete structures if used without scrutiny. In fairness contexts, such errors carry high stakes. False positives (e.g., unjustified bias paths) can trigger overcorrections, while false negatives obscure actual harms. We find that even underperforming methods can appear fairness-aware using naive metrics like $C_{\text{bias}}$, which may be inflated by erroneous causal paths. This highlights the need for path-level interpretability rather than summary statistics alone.
To mitigate these risks, our approach integrates LLMs as prior-informed agents within a selective AL loop. This reduces unnecessary queries, avoids early commitment errors, and improves efficiency over fixed-query strategies such as BFS. Like traditional DAG methods, our framework embraces uncertainty and treats CD as hypothesis generation.
While it is argued that LLMs may not perform causal reasoning per se, we treat them as semantic priors filtered through statistical signals and query history. This is analogous to expert elicitation—subject to bias, but useful when paired with empirical validation. As argued by (Imbens 2020) and (McCoy et al. 2023), epistemic uncertainty is intrinsic to CD. LLM-guided proposals can still surface plausible structures, especially in fairness-critical settings. Prior work has emphasized this role (Kıcıman et al. 2023); we build on it with a dynamic querying process that improves recovery while reducing false positives in bias pathways.
Our broader aim is to support interpretable, efficient, and socially aware fairness diagnostics, offering complementary insight to notions like demographic parity or counterfactual fairness. By uncovering how sensitive variables influence outcomes through direct and mediated paths, our method moves beyond outcome disparities toward causal insight. This aligns with calls for path-level transparency in AI audits from fairness, accountability, and legal communities. We see this work as a step toward embedding causal reasoning into policy evaluations, model and risk assessments, tools needed to build more trustworthy socio-technical systems.
This approach provides actionable insights for non-technical stakeholders. By surfacing fairness-relevant pathways, the method enables practitioners, compliance teams, or policy analysts to trace how sensitive attributes indirectly shape outcomes. These insights can guide interventions, from revising decision rules to informing accountability processes in hiring, lending, or regulatory oversight.